# Lean Finder: Semantic Search for Mathlib That Understands User Intents

### v1

@@ 1 Introduction
Advances in Lean and mathlib (De Moura et al., 2015; Moura & Ullrich, 2021) are turning mathematical discovery into a collaborative and verifiable research workflow. On this foundation, provers powered by large language models (LLMs) have sprinted ahead (AlphaProof & teams, 2024; Xin et al., 2024a; b; Ren et al., 2025; Lin et al., 2025a). Parallel progress in autoformalization has shifted from hand-written grammars to few-shot LLM translation, and large corpora generated through back-translating Lean code (“informalization”) have bootstrapped even stronger translators (Wu et al., 2022; Jiang et al., 2024; Lu et al., 2024).
Despite these advances, state-of-the-art LLMs still cannot solve math research problems. This underscores the continued need for substantial human efforts, which unfortunately remain slow and labor-intensive for two bottlenecks. First, locating the right theorem is frustrating: search tools are rudimentary, naming conventions are often inconsistent (Zulip, 2021b; 2020b), and high-quality Lean examples remain scarce. Moreover, as we will demonstrate in our experiments (Table LABEL:table:informal_user_header), identifying the correct theorem is challenging not only for humans, but also for state-of-the-art LLMs. Second, Lean’s syntax, grammar, and tactics incur a steep learning curve. Even veteran mathematicians and experienced programmers regularly report difficulties when writing Lean (Zulip, 2021a; 2020a). To facilitate mathematicians, semantic search is vital to Lean formalization and theorem proving. Recent search engines take informal statements or live proof states and retrieve mathlib4 lemmas or tactic suggestions, thereby aiding human Lean users (Gao et al., 2024a; b; Tao et al., 2025a; Shen et al., 2025; Ju & Dong, 2025; Asher, 2025). However, these search engines target machine translation rather than human use. They “informalize” statements from a supposedly neutral viewpoint, whereas real users bring inherent bias, typically seeking explanations from their own specific perspective.
@@
Consider the two queries below. The first is an informalization of a formal statement that current Lean search engines handle (Gao et al., 2024a; b; Ju & Dong, 2025; Asher, 2025):
@@
However, a human query [^1] We synthesized this “Query 2” (see Section 3.1), as we do not release any real user queries due to privacy concerns (Section M). in practice may look like:
@@ 2.1 Background: Search Engine for Lean and Mathlib
Lean (De Moura et al., 2015; Moura & Ullrich, 2021) is an interactive theorem prover based on dependent type theory, with a community-run library (mathlib4) containing over 230k theorems and 110k definitions. This offers far more “entry points” than typical programming platforms (Python exposes 89 built-in functions; C++ standard library has 105 headers). Finding the right lemma in this vast library is difficult: existing tools like #find, library_search, and Loogle [^2] https://loogle.lean-lang.org/ depend on exact names or goal states and often fail when naming conventions drift or examples are sparse. Recent LLM-based search engines (Gao et al., 2024a; b; Yang et al., 2023; Tao et al., 2025a; Shen et al., 2025; Ju & Dong, 2025; Asher, 2025; Zhu et al., 2025) embed informal queries or proof goals, but remain optimized for embedding similarity rather than the nuanced, shifting intents of mathematicians. In short, sheer library scale and rudimentary search together make locating lemmas a first-order pain point, even before one tackles Lean’s non-trivial syntax and tactic language.
@@ 3.1 User-centered Query Synthesis
Therefore, preparing a large fully annotated set of genuine user queries is unrealistic. Instead, we propose a novel reverse strategy to synthesize a large amount of diverse and realistic queries. The core idea is: based on different perspectives that mathematicians ask Lean/mathlib4-related questions (Section 3.1.1), we prompt LLM generation to simulate mathematicians’ intents (Section 3.1.2).
@@ 3.1.1 Semantic Analysis of Real Human Query
To achieve this semantic analysis, we utilize Lean Zulip Chat [^3] https://leanprover.zulipchat.com/, the primary discussion forum for Lean, as well as GitHub, as the main sources of real-world queries. We use the Zulip API to extract high-quality discussions from five active and thematically rich channels (new members, lean4, mathlib4, Is there code for X, metaprogramming/tactics), retaining up to the first five user messages, which typically capture the core query. We then prompt GPT-4o (Appendix N.2) to filter for questions answerable by Lean statements and paraphrase the main query articulated in the messages. This process yields 693 real user discussions from Zulip and Github [^4] Due to privacy concerns, we do not release either Zulip or GitHub collections..
@@
See Appendix N.3 and Appendix N.4 for the prompts. This process results in five distinct and semantically meaningful clusters of queries, as shown in Table 7, capturing a comprehensive range of intents from which mathematicians ask about Lean/mathlib4.
@@ 3.1.2 Query Synthesis with User Intents
We aim to instruct the LLM to generate plausible user queries that align with the specified cluster’s perspective while remaining grounded in the given formal content. To avoid forcing every statement into all intent clusters, we adopt a two-step procedure. First, we prompt GPT-4o to analyze the given Lean statement and determine which query clusters from Table 7 are applicable. This filtering step ensures that only semantically meaningful queries are generated, avoiding unnatural cases where a cluster is irrelevant to a statement. Second, for each selected cluster, we prompt GPT-4o with rich context information (e.g. formal/informal Lean statement, cluster information) to synthesize the queries. See Appendix N.5 for our prompt. This procedure yields a total of 582102 synthetic queries distributed across five distinct intent clusters, enabling fine-tuning of embedding models to better meet mathematicians’ needs on Lean search.
@@
3.2 The Largest Code Search Dataset for Lean Repositories
@@ 4 Experiments
Our experiments primarily compare Lean Finder with other Lean search engines across different input modalities, with particular emphasis on user-centered queries reflecting mathematicians’ intents. In Section 4.1, we evaluate retrieval performance across input modalities, followed by a user study on real queries in Section 4.2. Section 4.3 explores Lean Finder as a retrieval backbone for LLM-based provers, and Appendix J reports ablation studies on the impact of preference alignment and using synthetic queries in training.
@@ Results.
Lean Finder
64.2 88.9 93.3 0.75 54.4 84.4 91.4 0.68 82.7 97.0 97.7 0.89
GPT-4o (full name match)
14.8 20.9 22.6 13.6 21.5 23.3 39.7 42.9 44.1
GPT-4o (stem match)
21.1 28.4 30.0 17.8 27.4 30.5 48.2 52.6 54.1
@@
Lean Finder
24.6 56.8 67.9 0.40 8.3 30.1 40.0 0.19
GPT-4o (full name match)
7.4 13.3 15.0 4.5 8.5 9.9
GPT-4o (stem match)
10.1 17.3 19.7 6.4 11.5 13.6
@@ 5 Related Works
Prior efforts in code search (Husain et al., 2019; Wang et al., 2023), Lean retrieval (Gao et al., 2024a; b; Asher, 2025; Lean FRO, ; Morph Labs, ), and autoformalization (Jiang et al., 2023; Lin et al., 2024; Wang et al., 2024; Ying et al., 2024) have advanced retrieval and translation between natural language and formal mathematics, but remain limited by narrow input types, paraphrase-level alignment, or brittle translations. Our work complements these directions by introducing a user-centered, intent-aware retrieval framework that explicitly models mathematicians’ real query styles and integrates them into training, enabling Lean Finder to generalize across diverse modalities and substantially outperform existing systems on real user queries. Extended related work is provided in Appendix A.
@@ Appendix A Extended Related Works
Code search, aimed at retrieving the most semantically relevant code snippets from a codebase according to a specified natural language query, is a common activity that plays an important role in software development (Husain et al., 2019; Wang et al., 2023) and aligns closely with premise selection: both problems require retrieving a small set of code blocks that are helpful for downstream tasks. Early methods relied on lexical heuristics (Alama et al., 2014; Urban, 2004), while recent work leverages neural encoders (Irving et al., 2016; Wang & Deng, 2020; Mikuła et al., 2023; Yeh et al., 2023) and dense retrieval models (Karpukhin et al., 2020; Qu et al., 2020) by learning semantic embeddings. In formal domains like Lean, retrieval systems like Lean Search (Gao et al., 2024a; b) assist users by allowing them to search for relevant mathlib statements using natural language queries. These systems primarily function by matching the natural language translation of a formal statement to other formal statements in mathlib. While effective in some settings, this approach supports only narrow input types and treats informal queries as approximate paraphrases of formal code and focuses on surface-level alignment, without capturing the diverse intents behind real-world user questions or the broader context in which they arise. Lean Finder addresses this gap by introducing a user-centered, intent-aware retrieval framework for Lean that supports a broad range of input modalities and explicitly bridges the gap between the natural queries posed by mathematicians and the formal statements stored in Lean projects. By modeling real user intent and query diversity, Lean Finder achieves consistently strong retrieval performance across all modalities and demonstrates substantial gains over existing systems in a user study based on real Lean queries.
@@
Autoformalization, the automated translation of informal mathematics into formal proof-assistant code, has become a key strategy for addressing the scarcity of high-quality formal data. Early neural encoders for statement formalization (Wang et al., 2018) have been eclipsed by LLM-based pipelines that translate at scale with few-shot prompting or fine-tuning (Wu et al., 2022; 2024; Xin et al., 2024a; b). Systems such as DeepSeek-Prover, InternLM2.5-StepProver, and Goedel Prover series formalize vast Lean corpora for expert iteration (Xin et al., 2024a; Wu et al., 2024; Lin et al., 2025a; b), while others generate synthetic formal proofs without informal seeds (Wu et al., 2020; Wang & Deng, 2020; Poesia et al., 2024). Projects such as MMA, Lean-STaR, TheoremLlama, and Lean Workbook create natural-language/formal-language (NL–FL) pairs to mitigate data scarcity (Jiang et al., 2023; Lin et al., 2024; Wang et al., 2024; Ying et al., 2024); yet LLM brittleness still introduces subtle logical errors. Recent LLM tools embedded in user workflows (e.g., LeanDojo, LeanCopilot, LLM-Step) further streamline manual formalization by suggesting premises and proof tactics (Yang et al., 2023; Song et al., 2023; Welleck & Saha, 2023; Shen et al., 2025). Our work tackles the complementary challenge of autoformalization: synthesizing intent-rich, user-style queries from formal Lean statements to better serve practitioners. Unlike previous informalization approaches that simply translate formal statements into natural language, our pipeline explicitly models user intent and query phrasing. Injecting this intent-aware signal during fine-tuning yields significant retrieval improvements on real user queries, and bridges the gap between informal user needs and the formal knowledge encoded in Lean libraries.
@@ B.2 Summary of Our Dataset
For the input modality of synthesized user query, we show the clusters information and their corresponding examples in Table 7.
Intent Cluster Semantic Definition Examples
@@ Appendix C OpenAI API Configurations
Inspired by Herald Gao et al. (2024b), the prompt for informalization is shown in Section N.1.
@@
We employ GPT-4o to filter out user queries that cannot be addressed by any Lean 4 formal statement (Section 3.1.1). The temperature is set to 0.0 with a maximum of 500 output tokens. The prompt for query filtering is shown in Section N.2.
@@
  • Bootstrap Categorization. An initial set of representative user queries is clustered using a temperature of 1.0 and a token limit of 20000. The prompt used for initial cluster generation is shown in Section N.3.
  • Progressive Clustering. Remaining queries are progressively categorized with the same temperature and token settings. The prompt used for this part is shown in Section N.4.
@@
To synthesize diverse user queries (Section 3.1.2), we use the GPT-4o API with a temperature of 0.7 and a maximum output of 1000 tokens. The prompt used is shown in Section  N.5.
@@ I.2 Step-wise Theorem Proving Model
Appendix J Ablation Studies
@@
Lean Finder
64.2 88.9 93.3 0.75 54.4 84.4 91.4 0.68 82.7 97.0 97.7 0.89
w/o DPO
66.9 91.2 95.1 0.77 58.1 86.8 93.1 0.71 84.9 97.4 98.4 0.91
w/o SynQ+DPO
61.1 87.5 91.6 0.73 38.4 66.4 73.7 0.51 86.4 96.8 98.0 0.91
@@
Lean Finder
24.6 56.8 67.9 0.40 8.3 30.1 40.0 0.19
w/o DPO
27.7 62.3 74.1 0.44 8.8 29.5 39.9 0.19
w/o SynQ+DPO
28.5 62.5 73.7 0.44 8.7 28.4 41.0 0.18
Appendix K Module-wise Analysis
@@
Appendix L Lean Finder’s User Interface
@@
Appendix M Privacy and Data Release
@@
Appendix N Prompts
N.1 Prompt used for Lean statement informalization
N.2 Prompt used for user query filtering
N.3 Prompt for generating initial clusters of user queries
N.4 Prompt for progressive clustering of user queries
N.5 Prompt for generating user queries
Appendix O Datasheets for Datasets
@@
O.4 Preprocessing / Cleaning / Labeling
@@
Appendix P Misc.
URL to our web service, model, and benchmark. https://leanfinder.github.io We will continually update our system to support more Lean projects and a growing database of Lean statements.

### v2

@@ 1 Introduction
Advances in Lean and mathlib (De Moura et al., 2015; Moura and Ullrich, 2021) are turning mathematical discovery into a collaborative and verifiable research workflow. On this foundation, provers powered by large language models (LLMs) have sprinted ahead (AlphaProof and teams, 2024; Xin et al., 2024a; b; Ren et al., 2025; Lin et al., 2025a). Parallel progress in autoformalization has shifted from hand-written grammars to few-shot LLM translation, and large corpora generated through back-translating Lean code (“informalization”) have bootstrapped even stronger translators (Wu et al., 2022; Jiang et al., 2024; Lu et al., 2024).
Despite these advances, state-of-the-art LLMs still cannot solve math research problems. This underscores the continued need for substantial human efforts, which unfortunately remain slow and labor-intensive for two bottlenecks. First, locating the right theorem is frustrating: search tools are rudimentary, naming conventions are often inconsistent (Zulip, 2021b; 2020b), and high-quality Lean examples remain scarce. Moreover, as we will demonstrate in our experiments (Table LABEL:table:informal_user_header), identifying the correct theorem is challenging not only for humans, but also for state-of-the-art LLMs. Second, Lean’s syntax, grammar, and tactics incur a steep learning curve. Even veteran mathematicians and experienced programmers regularly report difficulties when writing Lean (Zulip, 2021a; 2020a). To facilitate mathematicians, semantic search is vital to Lean formalization and theorem proving. Recent search engines take informal statements or live proof states and retrieve mathlib4 lemmas or tactic suggestions, thereby aiding human Lean users (Gao et al., 2024a; b; Tao et al., 2025a; Shen et al., 2025; Ju and Dong, 2025; Asher, 2025). However, these search engines target machine translation rather than human use. They “informalize” statements from a supposedly neutral viewpoint, whereas real users bring inherent bias, typically seeking explanations from their own specific perspective.
@@
Consider the two queries below. The first is an informalization of a formal statement that current Lean search engines handle (Gao et al., 2024a; b; Ju and Dong, 2025; Asher, 2025):
@@
However, a human query [^1] We synthesized this “Query 2” (see Section 3.1), as we do not release any real user queries due to privacy concerns (Section N). in practice may look like:
@@ 2.1 Background: Search Engine for Lean and Mathlib
Lean (De Moura et al., 2015; Moura and Ullrich, 2021) is an interactive theorem prover based on dependent type theory, with a community-run library (mathlib4) containing over 230k theorems and 110k definitions. This offers far more “entry points” than typical programming platforms (Python exposes 89 built-in functions; C++ standard library has 105 headers). Finding the right lemma in this vast library is difficult: existing tools like #find, library_search, and Loogle [^2] https://loogle.lean-lang.org/ depend on exact names or goal states and often fail when naming conventions drift or examples are sparse. Recent LLM-based search engines (Gao et al., 2024a; b; Yang et al., 2023; Tao et al., 2025a; Shen et al., 2025; Ju and Dong, 2025; Asher, 2025; Zhu et al., 2025) embed informal queries or proof goals, but remain optimized for embedding similarity rather than the nuanced, shifting intents of mathematicians. In short, sheer library scale and rudimentary search together make locating lemmas a first-order pain point, even before one tackles Lean’s non-trivial syntax and tactic language.
@@ 3.1 User-centered Query Synthesis
Therefore, preparing a large fully annotated set of genuine user queries is unrealistic. Instead, we propose a novel reverse strategy to synthesize a large amount of diverse and realistic queries, as overviewed in Figure 3. The core idea is: based on different perspectives that mathematicians ask Lean/mathlib4-related questions (Section 3.1.1), we prompt LLM generation to simulate mathematicians’ intents (Section 3.1.2).
@@ 3.1.1 Semantic Analysis of Real Human Query
To achieve this semantic analysis, we utilize Lean Zulip Chat [^3] https://leanprover.zulipchat.com/, the primary discussion forum for Lean, as well as GitHub, as the main sources of real-world queries. We use the Zulip API to extract high-quality discussions from five active and thematically rich channels (new members, lean4, mathlib4, Is there code for X, metaprogramming/tactics), retaining up to the first five user messages, which typically capture the core query. We then prompt GPT-4o (Appendix O.2) to filter for questions answerable by Lean statements and paraphrase the main query articulated in the messages. This process yields 693 real user discussions from Zulip and Github [^4] Due to privacy concerns, we do not release either Zulip or GitHub collections..
@@
See Appendix O.3 and Appendix O.4 for the prompts. This process results in five distinct and semantically meaningful clusters of queries, as shown in Table 1, capturing a comprehensive range of intents from which mathematicians ask about Lean/mathlib4.
Intent Cluster Semantic Definition Examples
@@ 3.1.2 Query Synthesis with User Intents
We aim to instruct the LLM to generate plausible user queries that align with the specified cluster’s perspective while remaining grounded in the given formal content. To avoid forcing every statement into all intent clusters, we adopt a two-step procedure. First, we prompt GPT-4o to analyze the given Lean statement and determine which query clusters from Table 1 are applicable. This filtering step ensures that only semantically meaningful queries are generated, avoiding unnatural cases where a cluster is irrelevant to a statement. Second, for each selected cluster, we prompt GPT-4o with rich context information (e.g. formal/informal Lean statement, cluster information) to synthesize the queries. See Appendix O.5 for our prompt. This procedure yields a total of 582102 synthetic queries distributed across five distinct intent clusters, enabling fine-tuning of embedding models to better meet mathematicians’ needs on Lean search.
@@
3.2 The Largest Code Search Dataset for Lean Repositories
@@ 4 Experiments
Our experiments primarily compare Lean Finder with other Lean search engines across different input modalities, with particular emphasis on user-centered queries reflecting mathematicians’ intents. In Section 4.1, we evaluate retrieval performance across input modalities, followed by a user study on real queries in Section 4.2. Section 4.3 explores Lean Finder as a retrieval backbone for LLM-based provers, and Appendix K reports ablation studies on the impact of preference alignment and using synthetic queries in training. Appendix J reports comparisons with additional baselines such as Lean Explore (Asher, 2025) and Qwen3-Embedding-8B (Zhang et al., 2025).
@@ Results.
Lean Finder 64.2 88.9 93.3 0.75 54.4 84.4 91.4 0.68 82.7 97.0 97.7 0.89
GPT-4o (full name match) 14.8 20.9 22.6 13.6 21.5 23.3 39.7 42.9 44.1
GPT-4o (stem match) 21.1 28.4 30.0 17.8 27.4 30.5 48.2 52.6 54.1
@@
Lean Finder 24.6 56.8 67.9 0.40 8.3 30.1 40.0 0.19
GPT-4o (full name match) 7.4 13.3 15.0 4.5 8.5 9.9
GPT-4o (stem match) 10.1 17.3 19.7 6.4 11.5 13.6
@@ 5 Related Works
Prior efforts in code search (Husain et al., 2019; Wang et al., 2023), Lean retrieval (Gao et al., 2024a; b; Asher, 2025; Lean FRO, ; Morph Labs, ), and autoformalization (Jiang et al., 2023; Lin et al., 2024; Wang et al., 2024; Ying et al., 2024; Soroco et al., 2025) have advanced retrieval and translation between natural language and formal mathematics, but remain limited by narrow input types, paraphrase-level alignment, or brittle translations. Our work complements these directions by introducing a user-centered, intent-aware retrieval framework that explicitly models mathematicians’ real query styles and integrates them into training, enabling Lean Finder to generalize across diverse modalities and substantially outperform existing systems on real user queries. Extended related work is provided in Appendix A.
@@ 6 Conclusion
Acknowledgment
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0034682 and ASCR-ERCAP0031463.
@@ Appendix A Extended Related Works
Code search, aimed at retrieving the most semantically relevant code snippets from a codebase according to a specified natural language query, is a common activity that plays an important role in software development (Husain et al., 2019; Wang et al., 2023) and aligns closely with premise selection: both problems require retrieving a small set of code blocks that are helpful for downstream tasks. Early methods relied on lexical heuristics (Alama et al., 2014; Urban, 2004), while recent work leverages neural encoders (Irving et al., 2016; Wang and Deng, 2020; Mikuła et al., 2023; Yeh et al., 2023) and dense retrieval models (Karpukhin et al., 2020; Qu et al., 2020) by learning semantic embeddings. In formal domains like Lean, retrieval systems like Lean Search (Gao et al., 2024a; b) assist users by allowing them to search for relevant mathlib statements using natural language queries. These systems primarily function by matching the natural language translation of a formal statement to other formal statements in mathlib. While effective in some settings, this approach supports only narrow input types and treats informal queries as approximate paraphrases of formal code and focuses on surface-level alignment, without capturing the diverse intents behind real-world user questions or the broader context in which they arise. Lean Finder addresses this gap by introducing a user-centered, intent-aware retrieval framework for Lean that supports a broad range of input modalities and explicitly bridges the gap between the natural queries posed by mathematicians and the formal statements stored in Lean projects. By modeling real user intent and query diversity, Lean Finder achieves consistently strong retrieval performance across all modalities and demonstrates substantial gains over existing systems in a user study based on real Lean queries.
@@
Autoformalization, the automated translation of informal mathematics into formal proof-assistant code, has become a key strategy for addressing the scarcity of high-quality formal data. Early neural encoders for statement formalization (Wang et al., 2018) have been eclipsed by LLM-based pipelines that translate at scale with few-shot prompting or fine-tuning (Wu et al., 2022; 2024; Xin et al., 2024a; b; Soroco et al., 2025). Systems such as DeepSeek-Prover, InternLM2.5-StepProver, and Goedel Prover series formalize vast Lean corpora for expert iteration (Xin et al., 2024a; Wu et al., 2024; Lin et al., 2025a; b), while others generate synthetic formal proofs without informal seeds (Wu et al., 2020; Wang and Deng, 2020; Poesia et al., 2024). Projects such as MMA, Lean-STaR, TheoremLlama, and Lean Workbook create natural-language/formal-language (NL–FL) pairs to mitigate data scarcity (Jiang et al., 2023; Lin et al., 2024; Wang et al., 2024; Ying et al., 2024); yet LLM brittleness still introduces subtle logical errors. Recent LLM tools embedded in user workflows (e.g., LeanDojo, LeanCopilot, LLM-Step) further streamline manual formalization by suggesting premises and proof tactics (Yang et al., 2023; Song et al., 2023; Welleck and Saha, 2023; Shen et al., 2025). Our work tackles the complementary challenge of autoformalization: synthesizing intent-rich, user-style queries from formal Lean statements to better serve practitioners. Unlike previous informalization approaches that simply translate formal statements into natural language, our pipeline explicitly models user intent and query phrasing. Injecting this intent-aware signal during fine-tuning yields significant retrieval improvements on real user queries, and bridges the gap between informal user needs and the formal knowledge encoded in Lean libraries.
@@ Appendix C OpenAI API Configurations
Inspired by Herald Gao et al. (2024b), the prompt for informalization is shown in Section O.1.
@@
We employ GPT-4o to filter out user queries that cannot be addressed by any Lean 4 formal statement (Section 3.1.1). The temperature is set to 0.0 with a maximum of 500 output tokens. The prompt for query filtering is shown in Section O.2.
@@
  • Bootstrap Categorization. An initial set of representative user queries is clustered using a temperature of 1.0 and a token limit of 20000. The prompt used for initial cluster generation is shown in Section O.3.
  • Progressive Clustering. Remaining queries are progressively categorized with the same temperature and token settings. The prompt used for this part is shown in Section O.4.
@@
To synthesize diverse user queries (Section 3.1.2), we use the GPT-4o API with a temperature of 0.7 and a maximum output of 1000 tokens. The prompt used is shown in Section  O.5.
@@ I.2 Step-wise Theorem Proving Model
Appendix J Additional Experiments
We further compare Lean Finder with Lean Search that has augmentation mode enabled, Qwen3-Embedding-8B (Zhang et al., 2025), and Lean Explore (Asher, 2025) on their supported input types. For Lean Search with augmentation mode enabled, we first use its built-in query augmentation functionality to augment the query, then submit it to Lean Search for retrieval. Since Lean Search and Lean Explore doesn’t support proof state as input, we do not compare their retrieval performance when the input is of type proof state. We require the ground truth formal statement to be present in the respective static databases for both Lean Search and Lean Explore for fair comparison. Because this filtering process yields different test sets for each, we report their results separately.
Table 10: Retrieval performance comparison between Lean Finder, Lean Search with augmentation mode enabled, and Qwen3-Embedding-8B across three input modalities: informalized statement, synthetic user query, and augmented statement.The best results are presented in bold.
R@1 R@5 R@10 MRR R@1 R@5 R@10 MRR R@1 R@5 R@10 MRR
Lean Finder 64.2 88.9 93.3 0.75 54.4 84.4 91.4 0.68 82.7 97.0 97.7 0.89
Lean Search (augmentation mode) 31.9 77.4 84.8 0.44 30.0 72.7 83.3 0.43 45.4 86.0 93.0 0.58
Qwen3-Embedding-8B 57.7 82.5 89.2 0.69 41.2 73.8 82.8 0.55 76.2 93.4 95.1 0.83
Table 11: Comparison of Lean Finder with Qwen3-Embedding-8B on retrieval performance across two input modalities: augmented proof state and raw proof state. “Augmented proof state”: proof state augmented with natural language description of proofs (Section 3.2). “Raw proof state”: proof state without the augmentation. The best results are presented in bold.
Model Augmented Proof State Raw Proof State
R@1 R@5 R@10 MRR R@1 R@5 R@10 MRR
Lean Finder 24.6 56.8 67.9 0.40 8.3 30.1 40.0 0.19
Qwen3-Embedding-8B 8.0 26.1 34.7 0.17 5.4 21.1 27.4 0.13
Table 12: Retrieval performance comparison between Lean Finder and Lean Explore on the input types that Lean Explore support: informalized statement, synthetic user query, and augmented statement. The test set used is a cleaned version of the original test set to ensure the ground truth Lean statement is present in Lean Explore’s static database.The best results are presented in bold.
R@1 R@5 R@10 MRR R@1 R@5 R@10 MRR R@1 R@5 R@10 MRR
Lean Finder 65.7 89.0 93.2 0.76 57.3 85.4 91.2 0.69 86.8 96.7 98.0 0.91
Lean Explore 35.0 61.8 68.0 0.46 26.3 51.3 60.1 0.37 85.8 93.6 95.2 0.89
Lean Finder consistently surpasses all baselines, including specialized Lean search engines (such as Lean Explore) and the state-of-the-art Qwen3-Embedding-8B model, which is larger than Lean Finder by 1B parameters. This indicates Lean Finder’s strong performance and robustness across input types.
Appendix K Ablation Studies
@@
Lean Finder 64.2 88.9 93.3 0.75 54.4 84.4 91.4 0.68 82.7 97.0 97.7 0.89
w/o DPO 66.9 91.2 95.1 0.77 58.1 86.8 93.1 0.71 84.9 97.4 98.4 0.91
w/o SynQ+DPO 61.1 87.5 91.6 0.73 38.4 66.4 73.7 0.51 86.4 96.8 98.0 0.91
@@
Lean Finder 24.6 56.8 67.9 0.40 8.3 30.1 40.0 0.19
w/o DPO 27.7 62.3 74.1 0.44 8.8 29.5 39.9 0.19
w/o SynQ+DPO 28.5 62.5 73.7 0.44 8.7 28.4 41.0 0.18
Appendix L Module-wise Analysis
@@
Appendix M Lean Finder’s User Interface
@@
Appendix N Privacy and Data Release
@@
Appendix O Prompts
O.1 Prompt used for Lean statement informalization
O.2 Prompt used for user query filtering
O.3 Prompt for generating initial clusters of user queries
O.4 Prompt for progressive clustering of user queries
O.5 Prompt for generating user queries
Appendix P Datasheets for Datasets
@@
P.4 Preprocessing / Cleaning / Labeling
@@
Appendix Q Misc.
URL to our web service, model, and benchmark. We will continually update our system to support more Lean projects and a growing database of Lean statements.